# On decomposability and subdifferential of the tensor nuclear norm

### v1

@@ 1 Introduction
where $\mathbf{T}=\mathbf{U}\mathbf{D}\mathbf{V}^{\textnormal{T}}$ is a compact singular value decomposition (SVD) and $\|\bullet\|_{*}$ and $\|\bullet\|_{\sigma}$ denote the nuclear norm and the spectral norm, respectively. They play a critical role in many matrix optimization problems and statistical learning, such as compressed sensing [45, Corollary 5], matrix recovery and completion [12, 13], multivariate regression [45, Corollary 3], phase retrieval [57, Chapter 10.4], recommender systems [57, Example 10.2], and robust principal component analysis (PCA) [11]; see a recent book [57] and the references therein for more examples.
@@
With the rapid developments of data sciences in various fields, the recent decade has witnessed a significant increase in research activities in tensors, the higher-order generalization of matrices. In the same spirit to matrices, the tensor nuclear norm is widely recognized as a convex surrogate for the intrinsic complexity of a tensor [17] and has triggered many applications in statistical learning such as tensor completion [61], tensor regression [50], and tensor robust PCA [19]. However, the nice properties of the matrix nuclear norm, (1) and (2), do not carry over directly to the tensor nuclear norm. This has made many low-rank tensor optimization problems difficult and statistical implications on low-rank tensors unclear. In fact, the only known result on the decomposability of the tensor nuclear norm is the so-called weak decomposability [50, Lemma 1] that only works for third-order tensors. Although it has found applications in tensor regression [50], it does not completely address the entire issue. For the subdifferential of the tensor nuclear norm, only two limited inclusions, [61, Lemma 1] and [62, Theorem 1], have been proposed. While these inclusions have been applied to analyze the statistical performance of tensor completion [61, 62], they are still unsatisfactory from the perspective of tensor analysis since many obvious subgradients have been excluded.
@@
The main reason behind the full decomposability and the subdifferential characterization of the matrix nuclear norm is that every matrix admits an SVD. As a contrast, only a very special class of tensors admits SVDs. While this unfortunate fact is primarily responsible for the gaps of the properties, it indeed makes the problems subtler and more interesting. This paper aims to offer a more in-depth understanding of the decomposability and the subdifferential of the tensor nuclear norm.
@@
As the first main result, we find that the tensor nuclear norm can in fact be fully decomposable if we carefully choose the tensor subspaces in which $\mathbf{T}$ and $\mathbf{S}$ reside as those in (1). It also directly points out why only a weak decomposability is possible for the subspace considered in [50, Lemma 1]. Our full decomposability applies to tensors of an arbitrary order, leading to a very natural generalization of the matrix case in (1). Moreover, we optimize the subspaces that allow the full decomposability and determine the largest possible such subspace pairs. The study has also resulted a dual byproduct, the decomposability of the tensor spectral norm, running in parallel to the tensor nuclear norm.
@@
The full decomposability of the tensor nuclear norm offers a tool to study its subdifferential. As the next main result, we propose novel subdifferential inclusions of the tensor nuclear norm. They strictly enlarge the inclusion proposed in [62, Theorem 1] that is the only known subdifferential inclusion for tensors of an arbitrary order. In particular, we show that a full spectral ball, $\bigl\{\mathbf{X}:\|\mathbf{X}\bigr\|_{\sigma}\leq 1\}$, can be imposed in (2) rather than $\|\mathbf{X}\|_{\sigma}\leq\frac{2}{d(d-1)}$ imposed in [62, Theorem 1], where $d$ is the order of the tensor. Our study indicates that there is no universal way to explicitly characterize the subdifferential of the tensor nuclear norm as that of the matrix nuclear norm in (2), supported by various approximations of the subdifferential and several interesting examples. Moreover, we investigate subgradients of the tensor nuclear norm and derive various bounds for the inclusion and exclusion of the subdifferential in all relevant subspaces.
@@
We believe that these developments can be important to applications, no matter in theory or in practice. As a precursor, we propose an immediate application to analyze the statistical performance of the nuclear-norm-based tensor robust PCA that aims to recover a planted low-rank ground-truth tensor superposed by a sparse corruption. In the matrix case, the robust PCA is already a remarkable instance of modern compressed sensing and has found many interesting applications such as video surveillance, face recognition, and community detection; see, e.g., [11, 30, 59] and the references therein. A key component underlying its outstanding performance is the nuclear norm minimization that tends to reveal low-rank solutions. To the best of our knowledge, the statistical performance of the tensor robust PCA has only been established for third-order tensors [19]. Our result applies to tensors of an arbitrary order and exactly recovers the matrix case [11, Theorem 1.1] when $d=2$. Moreover, it enjoys some looser conditions for the exact recovery than those required in [19] when $d=3$. One interesting insight from our analysis suggests that the conditions for the exact recovery of the tensor robust PCA of every order are highly likely to be identical.
@@ 2.2 Tensor operations
A tensor $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$ has $d$ modes. Fixing entry indices of d1d-1 modes except mode $k$ results a vector in $\mathbb{R}^{n_{k}}$, called a mode-$k$ fiber. For matrices, mode-$1$ fibers are columns and mode-$2$ fibers are rows. The mode-$k$ matricization of $\mathbf{T}$, denoted by $\mathbf{T}_{(k)}\in\mathbb{R}^{n_{k}\times\prod_{i\neq k}n_{i}}$, is to arrange mode-$k$ fibers to be the columns of the resulting matrix. The mode-$k$ product between $\mathbf{T}$ and a matrix $\mathbf{X}\in\mathbb{R}^{n\times n_{k}}$, denoted by $\mathbf{T}\times_{k}\mathbf{X}\in\mathbb{R}^{n_{1}\times\dots\times n_{k-1}\times n\times n_{k+1}\times\dots\times n_{d}}$, changes every mode-$k$ fiber of $\mathbf{T}$, say $\mathbf{v}\in\mathbb{R}^{n_{k}}$, to $\mathbf{X}\mathbf{v}\in\mathbb{R}^{n}$, a mode-$k$ fiber of $\mathbf{T}\times_{k}\mathbf{X}$; in another word,
@@
$\left\|\mathbf{x}_{1}\otimes\mathbf{x}_{2}\otimes\dots\otimes\mathbf{x}_{d}\right\|_{2}=\prod_{k=1}^{d}\|\mathbf{x}_{d}\|_{2}.$
@@
$\mathbf{T}(\bullet,\mathbf{x}_{2},\mathbf{x}_{3},\dots,\mathbf{x}_{d})=\mathbf{T}\times_{2}\mathbf{x}_{2}\times_{3}\mathbf{x}_{3}\dots\times_{d}\mathbf{x}_{d}\in\mathbb{R}^{n_{1}}$
becomes a vector. Similarly, $\mathbf{T}(\bullet,\bullet,\mathbf{x}_{3},\mathbf{x}_{4},\dots,\mathbf{x}_{d})\in\mathbb{R}^{n_{1}\times n_{2}}$ is a matrix, and so on. For a thorough introduction to tensor operations, we refer interested readers to Kolda and Bader [32], Ballard and Kolda [1], and Nie [46, Chapter 11].
@@ 2.3 Tensor spectral norm and nuclear norm
see, e.g., [37, 49, 20] for more details. A decomposition $\sum_{i=1}^{r}\lambda_{i}\,\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}$ of $\mathbf{T}$ in (4) that attains $\|\mathbf{T}\|_{*}$ is called a nuclear decomposition. These two definitions reduce to the matrix spectral norm and nuclear norm when $d=2$. However, unlike their matrix counterparts, computing the tensor spectral norm [26] and nuclear norm [20] are both NP-hard when $d\geq 3$.
@@
and so on.
The tensor spectral norm is the dual norm of the tensor nuclear norm, and vise versa.
@@
A proof can be found in, e.g., [38, Lemma 21]. An optimal tensor $\mathbf{Z}$ of either of the above two optimization problems is called a dual certificate of $\mathbf{T}$. In addition, for any nuclear decomposition $\mathbf{T}=\sum_{i=1}^{r}\lambda_{i}\,\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}$ with $\lambda_{i}>0$ for $i\in[r]$ and any dual certificate $\mathbf{Z}$ with $\|\mathbf{Z}\|_{\sigma}=1$ and $\|\mathbf{T}\|_{*}=\langle\mathbf{T},\mathbf{Z}\rangle$, one always has $\left\langle\mathbf{Z},\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}\right\rangle=1$ for any $i\in[r]$; see [20, Lemma 4.1].
@@
for any tensor $\mathbf{T}$; see, e.g., [15, Proposition 4.2].
3 Tensors subspaces
The essential reason resulting the full decomposability of the tensor nuclear norm lies in properly chosen tensor subspaces. In this section, we introduce tensor subspaces and elaborate their notations as a key step to appreciate the main results. To provide a better picture of the tensor subspaces, we also discuss orthogonal transformations and projections for tensors. Finally, we prove key properties of the tensor spectral norm and nuclear norm over subspaces. These properties set a foundation for the main theoretical developments. The section is presented in a self-contained way for readers with minimal background knowledge.
@@ 3.2 Orthogonal transformations and projections
Orthogonal transformations preserve many matrix properties, in particular to this paper, keeping the spectral norm and nuclear norm unchanged. Given any matrix $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}}$, there is an orthogonal matrix $\mathbf{P}\in\mathbb{R}^{n_{1}\times n_{1}}$ such that the nonzero entries of $\mathbf{P}\mathbf{T}$ only appear in the first $r_{1}$ rows where $r_{1}=\dim\bigl(\operatorname{sp}_{1}(\mathbf{T}\bigr))$. There is also an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n_{2}\times n_{2}}$ such that the nonzero entries of $\mathbf{T}\mathbf{Q}$ only appear in the first $r_{2}$ columns where $r_{2}=\dim\bigl(\operatorname{sp}_{2}(\mathbf{T})\bigr)$. Its obvious that $r_{1}=r_{2}$, the rank of $\mathbf{T}$.
@@
In fact, the matrix SVD makes the above even better. Since $r_{1}=r_{2}$, SVD results the top-left $r_{1}\times r_{1}$ submatrix of $\mathbf{S}$ to be diagonal with all of the singular values. With that structure, the decomposability of the matrix nuclear norm is straightforward and the subdifferential of the matrix nuclear norm admits an explicit and complete representation; see (1) and (2). However, the diagonal structure does not hold for tensors in general, and this makes the results of decomposability and subdifferential of the tensor nuclear norm unsatisfactory.
@@
$\left(\bigotimes_{k=1}^{d}\operatorname{p}_{\mathbb{V}_{k}}\right)(\mathbf{T})=\mathbf{T}\times_{1}\mathbf{P}_{1}\times_{2}\mathbf{P}_{2}\dots\times_{d}\mathbf{P}_{d}.$
In another word, $\mathbf{T}\times_{k}\mathbf{P}_{k}$ projects all mode-$k$ fibers of $\mathbf{T}$ onto the subspace $\mathbb{V}_{k}$. Therefore, we can take $\bigotimes_{k=1}^{d}\operatorname{p}_{\mathbb{V}_{k}}$ as performing vector projections $d$ times, in any order of the modes $1,2,\dots,d$, as mode products allow swapping.

### v2

@@ 1 Introduction
where $\mathbf{T}=\mathbf{U}\mathbf{D}\mathbf{V}^{\textnormal{T}}$ is a compact singular value decomposition (SVD) and $\|\bullet\|_{*}$ and $\|\bullet\|_{\sigma}$ denote the nuclear norm and the spectral norm, respectively. They play a critical role in many matrix optimization problems and statistical learning, such as compressed sensing [50, Corollary 5], matrix recovery and completion [13, 14], multivariate regression [50, Corollary 3], phase retrieval [67, Chapter 10.4], recommender systems [67, Example 10.2], and robust principal component analysis (PCA) [12]; see a recent book [67] and the references therein for more examples.
@@
With the rapid developments of data sciences in various fields, the recent decade has witnessed a significant increase in research activities in tensors, the higher-order generalization of matrices. In the same spirit to matrices, the tensor nuclear norm is widely recognized as a convex surrogate for the intrinsic complexity of a tensor [20] and has triggered many applications in statistical learning such as tensor completion [71], tensor regression [55], and tensor robust PCA [22]. However, the nice properties of the matrix nuclear norm, (1) and (2), do not carry over directly to the tensor nuclear norm. This has made many low-rank tensor optimization problems difficult and statistical implications on low-rank tensors unclear. In fact, the only known result on the decomposability of the tensor nuclear norm is the so-called weak decomposability [55, Lemma 1] that only works for third-order tensors. Although it has found applications in tensor regression [55], it does not completely address the entire issue. For the subdifferential of the tensor nuclear norm, only two limited inclusions, [71, Lemma 1] and [72, Theorem 1], have been proposed. While these inclusions have been applied to analyze the statistical performance of tensor completion [71, 72], they are still unsatisfactory from the perspective of tensor analysis since many obvious subgradients have been excluded.
@@
The main reason behind the full decomposability and the subdifferential characterization of the matrix nuclear norm is that every matrix admits an SVD. As a contrast, only a very special class of tensors, such as orthogonally decomposable tensors [73, 60, 57, 27], admits SVDs [17, Definition 4.1]. While this unfortunate fact is primarily responsible for the gaps of the properties, it indeed makes the problems subtler and more interesting. This paper aims to offer a more in-depth understanding of the decomposability and the subdifferential of the tensor nuclear norm.
@@
As the first main result, we find that the tensor nuclear norm can in fact be fully decomposable if we carefully choose the tensor subspaces in which $\mathbf{T}$ and $\mathbf{S}$ reside as those in (1). It also directly points out why only a weak decomposability is possible for the subspace considered in [55, Lemma 1]. Our full decomposability applies to tensors of an arbitrary order, leading to a very natural generalization of the matrix case in (1). Moreover, we optimize the subspaces that allow the full decomposability and determine the largest possible such subspace pairs. The study has also resulted in a dual byproduct, the decomposability of the tensor spectral norm, running in parallel to the tensor nuclear norm.
@@
The full decomposability of the tensor nuclear norm offers a tool to study its subdifferential. As the next main result, we propose novel subdifferential inclusions of the tensor nuclear norm. They strictly enlarge the inclusion proposed in [72, Theorem 1] that is the only known subdifferential inclusion for tensors of an arbitrary order. In particular, we show that a full spectral ball, $\bigl\{\mathbf{X}:\|\mathbf{X}\bigr\|_{\sigma}\leq 1\}$, can be imposed in (2) rather than $\|\mathbf{X}\|_{\sigma}\leq\frac{2}{d(d-1)}$ imposed in [72, Theorem 1], where $d$ is the order of the tensor. Our study indicates that there is no universal way to explicitly characterize the subdifferential of the tensor nuclear norm as that of the matrix nuclear norm in (2), supported by various approximations of the subdifferential and several interesting examples. Moreover, we investigate subgradients of the tensor nuclear norm and derive various bounds for the inclusion and exclusion of the subdifferential in all relevant subspaces.
@@
We believe that these developments can be important to applications, no matter in theory or in practice. As a precursor, we propose an immediate application to analyze the statistical performance of the nuclear-norm-based tensor robust PCA that aims to recover a planted low-rank ground-truth tensor superposed by a sparse corruption. In the matrix case, the robust PCA is already a remarkable instance of modern compressed sensing and has found many interesting applications such as video surveillance, face recognition, and community detection; see, e.g., [12, 34, 69] and the references therein. A key component underlying its outstanding performance is the nuclear norm minimization that tends to reveal low-rank solutions. To the best of our knowledge, the statistical performance of the tensor robust PCA has only been established for third-order tensors [22]. Our result applies to tensors of an arbitrary order and exactly recovers the matrix case [12, Theorem 1.1] when $d=2$. Moreover, it enjoys some looser conditions for the exact recovery than those required in [22] when $d=3$. One interesting insight from our analysis suggests that the conditions for the exact recovery of the tensor robust PCA of every order are highly likely to be identical.
@@ 2.2 Tensor operations
A tensor $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$ has $d$ modes. Fixing entry indices of d1d-1 modes except mode $k$ results in a vector in $\mathbb{R}^{n_{k}}$, called a mode-$k$ fiber. For matrices, mode-$1$ fibers are columns and mode-$2$ fibers are rows. The mode-$k$ matricization of $\mathbf{T}$, denoted by $\mathbf{T}_{(k)}\in\mathbb{R}^{n_{k}\times\prod_{i\neq k}n_{i}}$, is to arrange mode-$k$ fibers to be the columns of the resulting matrix. The mode-$k$ product between $\mathbf{T}$ and a matrix $\mathbf{X}\in\mathbb{R}^{n\times n_{k}}$, denoted by $\mathbf{T}\times_{k}\mathbf{X}\in\mathbb{R}^{n_{1}\times\dots\times n_{k-1}\times n\times n_{k+1}\times\dots\times n_{d}}$, changes every mode-$k$ fiber of $\mathbf{T}$, say $\mathbf{v}\in\mathbb{R}^{n_{k}}$, to $\mathbf{X}\mathbf{v}\in\mathbb{R}^{n}$, a mode-$k$ fiber of $\mathbf{T}\times_{k}\mathbf{X}$; in other words,
@@
$\left\|\mathbf{x}_{1}\otimes\mathbf{x}_{2}\otimes\dots\otimes\mathbf{x}_{d}\right\|_{2}=\prod_{k=1}^{d}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\definecolor{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\definecolor{XC@mixtmp}{rgb}{0,0,0}\definecolor{XC@mixtmp}{rgb}{0,0,0}\mathbf{x}_{k}}\|_{2}.$
@@
$\mathbf{T}(\bullet,\mathbf{x}_{2},\mathbf{x}_{3},\dots,\mathbf{x}_{d})=\mathbf{T}\times_{2}\mathbf{x}_{2}\times_{3}\mathbf{x}_{3}\dots\times_{d}\mathbf{x}_{d}\in\mathbb{R}^{n_{1}}$ (3)
becomes a vector. Similarly, $\mathbf{T}(\bullet,\bullet,\mathbf{x}_{3},\mathbf{x}_{4},\dots,\mathbf{x}_{d})\in\mathbb{R}^{n_{1}\times n_{2}}$ is a matrix, and so on. For a thorough introduction to tensor operations, we refer interested readers to Kolda and Bader [36], Ballard and Kolda [1], and Nie [51, Chapter 11].
@@ 2.3 Tensor spectral norm and nuclear norm
see, e.g., [42, 54, 23] for more details. A decomposition $\sum_{i=1}^{r}\lambda_{i}\,\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}$ of $\mathbf{T}$ in (5) is commonly known as a CANDECOMP/PARAFAC (CP) decomposition in the literature; see, e.g., [36, Section 3]. A CP decomposition of $\mathbf{T}$ that attains $\|\mathbf{T}\|_{*}$ is called a nuclear decomposition. These two definitions, (4) and (5), reduce to the matrix spectral norm and nuclear norm when $d=2$. However, unlike their matrix counterparts, computing the tensor spectral norm [30] and nuclear norm [23] are both NP-hard when $d\geq 3$.
@@
and so on. In particular, the second equality follows immediately from (3) and the Cauchy-Schwarz inequality. The third equality follows from the fact that $\mathbf{T}(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{d-2},\bullet,\bullet)=\mathbf{T}\times_{1}\mathbf{x}_{1}\times_{2}\mathbf{x}_{2}\dots\times_{d-2}\mathbf{x}_{d-2}\in\mathbb{R}^{n_{d-1}\times n_{d}}$ and the definition of matrix spectral norm, i.e., (4) for $d=2$. Similar reasoning applies to the last equality.
The tensor spectral norm is the dual norm of the tensor nuclear norm, and vice versa.
@@
A proof can be found in, e.g., [43, Lemma 21]. An optimal tensor $\mathbf{Z}$ of either of the above two optimization problems is called a dual certificate of $\mathbf{T}$. In addition, for any nuclear decomposition $\mathbf{T}=\sum_{i=1}^{r}\lambda_{i}\,\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}$ with $\lambda_{i}>0$ for $i\in[r]$ and any dual certificate $\mathbf{Z}$ with $\|\mathbf{Z}\|_{\sigma}=1$ and $\|\mathbf{T}\|_{*}=\langle\mathbf{T},\mathbf{Z}\rangle$, one always has $\left\langle\mathbf{Z},\mathbf{x}_{1}^{i}\otimes\mathbf{x}_{2}^{i}\otimes\dots\otimes\mathbf{x}_{d}^{i}\right\rangle=1$ for any $i\in[r]$; see [23, Lemma 4.1].
@@
for any tensor $\mathbf{T}$; see, e.g., [16, Proposition 4.2].
3 Tensor subspaces
The essential reason resulting in the full decomposability of the tensor nuclear norm lies in properly chosen tensor subspaces. In this section, we introduce tensor subspaces and elaborate their notations as a key step to appreciate the main results. To provide a better picture of the tensor subspaces, we also discuss orthogonal transformations and projections for tensors. Finally, we prove key properties of the tensor spectral norm and nuclear norm over subspaces. These properties set a foundation for the main theoretical developments. The section is presented in a self-contained way for readers with minimal background knowledge.
@@ 3.2 Orthogonal transformations and projections
Orthogonal transformations preserve many matrix properties, in particular to this paper, keeping the spectral norm and nuclear norm unchanged. Given any matrix $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}}$, there is an orthogonal matrix $\mathbf{P}\in\mathbb{R}^{n_{1}\times n_{1}}$ such that the nonzero entries of $\mathbf{P}\mathbf{T}$ only appear in the first $r_{1}$ rows where $r_{1}=\dim\bigl(\operatorname{sp}_{1}(\mathbf{T}\bigr))$. There is also an orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{n_{2}\times n_{2}}$ such that the nonzero entries of $\mathbf{T}\mathbf{Q}$ only appear in the first $r_{2}$ columns where $r_{2}=\dim\bigl(\operatorname{sp}_{2}(\mathbf{T})\bigr)$. It is obvious that $r_{1}=r_{2}$, the rank of $\mathbf{T}$.
@@
In fact, the matrix SVD makes the above even better. Since $r_{1}=r_{2}$, SVD results in the top-left $r_{1}\times r_{1}$ submatrix of $\mathbf{S}$ to be diagonal with all of the singular values. With that structure, the decomposability of the matrix nuclear norm is straightforward and the subdifferential of the matrix nuclear norm admits an explicit and complete representation; see (1) and (2). However, the diagonal structure does not hold for tensors in general, and this makes the results of decomposability and subdifferential of the tensor nuclear norm unsatisfactory.
@@
$\left(\bigotimes_{k=1}^{d}\operatorname{p}_{\mathbb{V}_{k}}\right)(\mathbf{T})=\mathbf{T}\times_{1}\mathbf{P}_{1}\times_{2}\mathbf{P}_{2}\dots\times_{d}\mathbf{P}_{d},$
known as the multilinear orthogonal projection; see, e.g., [19, Section 2.6] and [65, Section 3], as well as the discussions therein. In other words, $\mathbf{T}\times_{k}\mathbf{P}_{k}$ projects all mode-$k$ fibers of $\mathbf{T}$ onto the subspace $\mathbb{V}_{k}$. Therefore, we can take $\bigotimes_{k=1}^{d}\operatorname{p}_{\mathbb{V}_{k}}$ as performing vector projections $d$ times, in any order of the modes $1,2,\dots,d$, as mode products allow swapping.